### v2

@@ 1 Introduction
For the benefit of the readers, in the following subsections we will first give a self-contained mathematical description of the inverse problem at hand and survey some related literature. In the subsequent sections, we will describe our main theoretical results, give details about the proposed architecture and support our findings with numerical experiments. We begin now by summarizing our main contributions.
1.1 Contribution
Our contributions can be summarized in the following points:
  • 1.
    In this work, we propose a DeepONet architecture for solving the EIT inverse problem and provide a rigorous mathematical justification for using such an architecture.
  • 2.
    Various approximation-theoretic results for operator networks which seek to learn operators between infinite-dimensional function spaces have appeared in the recent past. We extend those results in order to make them suitable for the EIT inverse problem by proving a Universal Approximation Theorem type result for the case when the input space lies in the class of Hilbert-Schmidt operators (related to the current-to-voltage map) and the output space is a function space (corresponding to the conductivity function).
  • 3.
    We show that our proposed method achieves good reconstructions and show that it outperforms the state-of-the-art IRGN baseline.
@@ 1.2 Mathematical preliminaries and background of the EIT problem
$\nabla\cdot(\gamma\nabla u)$ $=0\quad\mbox{ in }\Omega;$ (1)
$\begin{split}\gamma\partial_{\nu}u|_{\partial\Omega}&=g\mbox{ on }\partial\Omega\quad(\text{or more generally }\Sigma)\\ &=0\mbox{ else.}\end{split}$ (2)
@@
$\Lambda_{\gamma}:L^{2}_{\diamond}(\partial\Omega)\to L^{2}_{\diamond}(\partial\Omega),\quad g\mapsto u|_{\partial\Omega}.$ (3)
To emphasize, $u|_{\partial\Omega}$ refer to the voltage measurements on the the boundary $\partial\Omega$. Additionally, here and below we will use $\Lambda_{1}$ to denote the background N-t-D operator corresponding to the background conductivity. The abstract model as described in this article is also called the Calderón problem and has been studied by many authors since it was first described in [9]. We would like to mention here, that in the abstract mathematical formulation, typically the Dirichlet-to-Neumann (D-t-N) map is considered instead of the N-t-D map as is considered in this work. However, in practical settings such as in EIT, the data that is generated is in the form of current-to-voltage maps which is best captured by the N-t-D maps. A more practical formulation of the EIT problem is given by the Complete Electrode Model (CEM), [17]. As before we let $\gamma\in L^{\infty}_{+}(\Omega)$. Let there be $M$ open, connected and mutually disjoint electrodes, $E_{m}\subset\partial\Omega$, $m=\{1,\cdots,M\}$ having the same contact impedance $z>0$. Assuming that the injected currents $I_{m}$ on the electrodes satisfy the Kirchoff condition, i.e. $\Sigma_{m=1}^{M}I_{m}=0$, the resulting electric potential $(u,U)\in H^{1}(\Omega)\times\mathbb{R}^{M}$ satisfies (1) along with the following conditions on the boundary:
$\gamma\partial_{\nu}u$ $=0;\quad\text{on }\partial\Omega\setminus\bigcup E_{m}$ (4)
$\int_{E_{m}}\gamma\partial_{\nu}udS$ $=I_{m}\quad\text{on }E_{m},m=1,\cdots,M$ (5)
$u+z\sigma\partial_{\nu}u$ $=U_{m}\quad\text{on }E_{m},m=\{1,\cdots,M\}$ (6)
where the vector $U=(U_{1},\cdots,U_{M})$ represents voltage measurement on the electrodes satisfying a ground condition $\Sigma_{m=1}^{M}U_{m}=0$. We can thus define the $M-$electrode current to voltage operator, $R_{M}(\gamma):\mathbb{R}^{M}_{\diamond}\to\mathbb{R}^{M}_{\diamond}:I:=(I_{1},\cdots,I_{M})\mapsto(U_{1},\cdots,U_{M})=U$ and $U=R_{M}(\gamma)I$. For later use, we also define the ‘shifted’ CEM operator, $\tilde{R}_{M}(\gamma):={R}_{M}(\gamma)-{R}_{M}(1)$ where ${R}_{M}(1)$ is the background CEM current-to-voltage operator. Since the background operator $\tilde{R}_{M}(1)$ is known, $\tilde{R}_{M}(\gamma)$ is completely determined by $R_{M}(\gamma)$ and vice-versa. Before we define the inverse problem at hand, we will first define the set of admissible conductivities considered in this work. We recall the following definition from [27, Definition 2.2].
Definition 1.
@@
Given such a finite dimesnional subset $\mathcal{F}$ and two numbers $m,M$ we denote the space of admissible conductivities $\Gamma$ as:
$\Gamma$ =$=\{\gamma\in\mathcal{F}:M>\gamma(x)\geq m>0,\gamma(x)=1\text{ in }\Omega\setminus\Omega^{\prime}\}$ (7)
@@ 2 The Learning Problem for EIT
Our goal in this article is to propose a DeepONet based neural-network architecture for learning to invert Neumann-to-Dirichlet operator relevant to the problem of EIT. To formally set up the learning problem, let us consider the shifted abstract N-t-D operator $\tilde{\Lambda}_{\gamma}={\Lambda}_{\gamma}-\Lambda_{1}$, where $\Lambda_{1}$ is the background N-t-D operator. Analogous to the CEM case, $\tilde{\Lambda}_{\gamma}$ is completely determined by $\Lambda_{\gamma}$ and vice-versa. Now we list some mapping properties of the operator $\tilde{\Lambda}_{\gamma}$ which will be crucial in our analysis. First of all, note that $\tilde{\Lambda}_{\gamma}$ is a smoothing opertor, see e.g. [30, Theorem A.3] and [20, (2.5)]. In particular the operator norm of $\tilde{\Lambda}_{\gamma}$ is bounded, i.e., $\lvert\lvert\tilde{\Lambda}_{\gamma}\rvert\rvert_{\mathcal{L}(H^{s}\to H^{t})}\leq C$ for all $s,t\in\mathbb{R}$. From [1, Lemma 18], we know that $\tilde{\Lambda}_{\gamma}\in\mathbb{H}_{r}$ for any $r$. Here $\mathbb{H}_{r}$ denotes a space of Hilbert-Schmidt operators, see [1, eq.10] which is a Hilbert-space in its own right. Consider the set $D_{\Lambda}=\{\tilde{\Lambda}_{\gamma}:\gamma\in\Gamma\}$ of N-t-D operators corresponding to admissible conductivities. We show in 3 that $D_{\Lambda}$ is a compact set of $\mathbb{H}_{r}$.
@@
Formally, the problem is to ‘learn’ the map $G^{\text{true}}:D_{\Lambda}\to\Gamma$. More particularly, we want to design a deep neural network (DeepONet) to approximate the map, $G^{\text{true}}$. For learning operators between infinite dimensional spaces, DeepONets were introduced in [50, 43]. These generalized the (shallow-net) architecture proposed in [12, 11]. Theoretical guarantees for DeepONets were given in [43] where the neural network tries to learn operators between two Hilbert spaces. This theory was further generalized in [39, 42] and an architecture called Fourier Neural Operators was introduced for learning operators between Banach spaces having a certain approximation property. With this background in mind, and to set the stage for formulating our learning problem as learning an operator between two Hilbert Spaces, let us introduce a (regular)‘link function’ $\Phi:\mathbb{R}\to[m,\infty)$ such that any conductivity $\gamma\in\Gamma$ is given by $\gamma=\Phi\circ f$ for some $f\in L^{\infty}(\Omega)\cap L^{2}{(\Omega)}:=\tilde{\Gamma}$. An example of such a link function is given in [55, Example 8]. Now we will modify our operator learning problem. For this we first define a map, $\mathcal{G}^{\dagger}:D_{\Lambda}\to\tilde{\Gamma}$ where $\mathcal{G}^{\dagger}=\phi^{-1}\circ G^{\text{true}}$. Thus knowing $\mathcal{G}^{\dagger}$ completely determines $G^{\text{true}}$. Next we extend this map: Let $\mathcal{G}^{\dagger}:D_{\Lambda}\to L^{\infty}(\Omega)\cap L^{2}(\Omega)$ be the map as above, then there exists an extension $\mathcal{G}^{\dagger}_{\text{ext}}:\mathbb{H}_{r}\to L^{\infty}(\Omega)\cap L^{2}(\Omega)$ such that $\mathcal{G}^{\dagger}_{\text{ext}}|_{D_{\Lambda}}=\mathcal{G}^{\dagger}$. The extension $\mathcal{G}^{\dagger}_{\text{ext}}$ exists by Lemma 4. The map $\mathcal{G}^{\dagger}_{\text{ext}}$ will be approximated by a deep neural operator called DeepONet. Essentially, a neural operator is a parametric map between the input space (e.g. space of N-t-D operators) and the output space (e.g. space of admissible conductivities) that can be interpreted as a composition of three maps, an encoder, an approximator and a reconstructor, see [43] and Figure 1. As such an upper bound on the error in approximating the true operator by a neural operator can be split into upper bounds on the encoder error, the approximator error and the reconstructor error correspondingly. Now we recall the DeepONet as proposed in [43, 50] and adapt the definitions as applicable to the case of EIT. To that end, we define the following operators while keeping the terminology the same as in [43]:
@@
    $\mathcal{E}_{M}:=\mathcal{E}:D_{\Lambda}\subset\mathbb{H}_{r}\to\mathbb{R}^{M}\times\mathbb{R}^{M};\quad\tilde{\Lambda}_{\gamma}\mapsto\tilde{R}_{M}(\gamma)$ (8)
    The encoder can be thought of as ‘sensor evaluations’ in the spirit of [43]. Whenever it is clear from context, we will denote $\mathcal{E}_{M}$ by $\mathcal{E}$. As shown in subsection 7.1, the map $\mathcal{E}$ is continuous and can be extended continuously on $\mathbb{H}_{r}$. With a slight abuse of notation, we call this extended map $\mathcal{E}$ too.
  • 2.
    Decoder: While the encoder is a projection of an infinite-dimensional object (i.e. $\tilde{\Lambda}_{\gamma}\in\mathbb{H}_{r}$) into a finite-dimensional space, a decoder lifts back a finite dimensional object into its corresponding representation in the space of Hilbert-Schmidt operators. This lifting is done using the construction given in [20, section 6] and results in an operator ${\Lambda}^{M}_{\gamma}$, see [20, eqn. 6.8]. Thus, we have:
    $\mathcal{D}_{M}:=\mathcal{D}:\mathbb{R}^{M}\times\mathbb{R}^{M}\to\mathbb{H}_{r};\quad\tilde{R}_{M}(\gamma)\mapsto\tilde{\Lambda}^{M}_{\gamma}.$ (9)
@@
    $\mathcal{A}:\mathbb{R}^{M}\times\mathbb{R}^{M}\to\mathbb{R}^{P};\quad\tilde{R}_{M}(\gamma)\mapsto\{\mathcal{A}_{k}\}_{k=1}^{P}$ (10)
@@
    $\mathbf{\tau}:\mathbb{R}^{d}\to\mathbb{R}^{P+1},\quad y=(y_{1},\cdots,y_{d})\mapsto\{\tau_{k}\}_{k=0}^{P}$ (11)
@@
    $\mathcal{R}_{P}:=\mathcal{R}:\mathbb{R}^{P}\to L^{2}(\Omega),\quad\mathcal{R}(\alpha_{k}):=\tau_{0}(y)+\Sigma_{k=1}^{P}\alpha_{k}\tau_{k}(y).$ (12)
  • 5.
@@
    $\mathcal{P}_{P}:=\mathcal{P}:L^{2}(\Omega)\to\mathbb{R}^{p},\quad\mathcal{P}(f)=\left(\langle f-\tau_{0},\tau_{1}^{*}\rangle,\cdots,\langle f-\tau_{0},\tau_{P}^{*}\rangle\right)$ (13)

### v3

@@ Abstract
keywords:
Electrical Impedance Tomography, DeepONets, Calderon’s Problem
{MSCcodes}
44A12, 65R10, 94A12
@@ 1 Introduction
For the benefit of the readers, in the following subsections we will first give a self-contained mathematical description of the inverse problem at hand and survey some related literature. In the subsequent sections, we will describe our main theoretical results, give details about the proposed architecture and support our findings with numerical experiments.
@@ 1.1 Mathematical preliminaries and background of the EIT problem
(1) $\nabla\cdot(\gamma\nabla u)$ $=0\quad\mbox{ in }\Omega;$
(2) $\begin{split}\gamma\partial_{\nu}u|_{\partial\Omega}&=g\mbox{ on }\partial\Omega\quad(\text{or more generally }\Sigma)\\ &=0\mbox{ else.}\end{split}$
@@
(3) $\Lambda_{\gamma}:L^{2}_{\diamond}(\partial\Omega)\to L^{2}_{\diamond}(\partial\Omega),\quad g\mapsto u|_{\partial\Omega}.$
To emphasize, $u|_{\partial\Omega}$ refer to the voltage measurements on the the boundary $\partial\Omega$. Additionally, here and below we will use $\Lambda_{1}$ to denote the background N-t-D operator corresponding to the background conductivity. We would like to mention here, that in the abstract mathematical formulation, typically the Dirichlet-to-Neumann (D-t-N) map is considered instead of the N-t-D map as is considered in this work. However, in practical settings such as in EIT, the data that is generated is in the form of current-to-voltage maps which is best captured by the N-t-D maps. A more practical formulation of the EIT problem is given by the Complete Electrode Model (CEM), [18]. As before we let $\gamma\in L^{\infty}_{+}(\Omega)$. Let there be $M$ open, connected and mutually disjoint electrodes, $E_{m}\subset\partial\Omega$, $m=\{1,\cdots,M\}$ having the same contact impedance $z>0$. Assuming that the injected currents $I_{m}$ on the electrodes satisfy the Kirchoff condition, i.e. $\Sigma_{m=1}^{M}I_{m}=0$, the resulting electric potential $(u,U)\in H^{1}(\Omega)\times\mathbb{R}^{M}$ satisfies (1) along with the following conditions on the boundary:
(4) $\gamma\partial_{\nu}u$ $=0;\quad\text{on }\partial\Omega\setminus\bigcup E_{m}$
(5) $\int_{E_{m}}\gamma\partial_{\nu}udS$ $=I_{m}\quad\text{on }E_{m},m=1,\cdots,M$
(6) $u+z\sigma\partial_{\nu}u$ $=U_{m}\quad\text{on }E_{m},m=\{1,\cdots,M\}$
where the vector $U=(U_{1},\cdots,U_{M})$ represents voltage measurement on the electrodes satisfying a ground condition $\Sigma_{m=1}^{M}U_{m}=0$. We can thus define the $M-$electrode current to voltage operator, $R_{M}(\gamma):\mathbb{R}^{M}_{\diamond}\to\mathbb{R}^{M}_{\diamond}:I:=(I_{1},\cdots,I_{M})\mapsto(U_{1},\cdots,U_{M})=U$ and $U=R_{M}(\gamma)I$. For later use, we also define the ‘shifted’ CEM operator, $\tilde{R}_{M}(\gamma):={R}_{M}(\gamma)-{R}_{M}(1)$ where ${R}_{M}(1)$ is the background CEM current-to-voltage operator. Since the background operator $\tilde{R}_{M}(1)$ is known, $\tilde{R}_{M}(\gamma)$ is completely determined by $R_{M}(\gamma)$ and vice-versa. Before we define the inverse problem at hand, we will first define the set of admissible conductivities considered in this work. We recall the following definition from [28, Definition 2.2].
Definition 1.1.
@@
Given such a finite dimensional subset $\mathcal{F}$ and two positive numbers $m,M$ we denote the space of admissible conductivities $\Gamma$ as:
(7) $\Gamma$ =$=\{\gamma\in\mathcal{F}:M>\gamma(x)\geq m>0,\gamma(x)=1\text{ in }\Omega\setminus\Omega^{\prime}\}$
@@ 2 The Learning Problem for EIT
Our goal in this article is to propose a DeepONet based neural-network architecture for learning to invert Neumann-to-Dirichlet operator relevant to the problem of EIT. To formally set up the learning problem, let us consider the shifted abstract N-t-D operator $\tilde{\Lambda}_{\gamma}={\Lambda}_{\gamma}-\Lambda_{1}$, where $\Lambda_{1}$ is the background N-t-D operator. Analogous to the CEM case, $\tilde{\Lambda}_{\gamma}$ is completely determined by $\Lambda_{\gamma}$ and vice-versa. Now we list some mapping properties of the operator $\tilde{\Lambda}_{\gamma}$ which will be crucial in our analysis. First of all, note that $\tilde{\Lambda}_{\gamma}$ is a smoothing operator, see e.g. [31, Theorem A.3] and [21, (2.5)]. In particular the operator norm of $\tilde{\Lambda}_{\gamma}$ is bounded, i.e., $\lvert\lvert\tilde{\Lambda}_{\gamma}\rvert\rvert_{\mathcal{L}(H^{s}\to H^{t})}\leq C$ for all $s,t\in\mathbb{R}$. It can be shown that the operator $\tilde{\Lambda}_{\gamma}$ belongs to the class of Hilbert-Schmidt operators. For readers’ convenience we recall the definition of a Hilbert-Schmidt operator from [1] here. Consider an orthonormal basis of $L^{2}(\partial\Omega)$ given by $\{\phi_{k}:=\phi_{k}^{(0)}\}_{k=0}^{\infty}$ consisting of real-valued eigenfunctions of the Laplace–Beltrami operator on the compact manifold $\partial\Omega$. By removing the constant function $\phi_{0}$ from the above collection, we obtain bases for $L^{2}(\partial\Omega)/\mathbb{C}\quad\text{and}\quad L^{2}_{\diamond}(\partial\Omega)=H^{0}_{\diamond}(\partial\Omega).$ Now, by an appropriately rescaling these basis functions we can also get orthonormal bases $\{\phi_{k}^{(r)}\}_{k=1}^{\infty}$ of all $H^{r}(\partial\Omega)/\mathbb{C}$ and $H^{r}_{\diamond}(\partial\Omega)$ spaces, $r\in\mathbb{R}$. For $j,k\in\mathbb{N}$, denote the tensor product operator, $b_{jk}^{(r)}:H^{r}(\partial\Omega)\to L^{2}(\partial\Omega)$
$b_{jk}^{(r)}(f)=\phi_{j}^{(r)}\otimes\phi_{k}^{(0)}(f):=\langle f,\phi_{j}^{(r)}\rangle_{H^{r}(\partial D)}\phi_{k}^{(0)},\quad f\in H^{r}(\partial\Omega),$
and define the space of linear operators
$\mathbb{H}_{r}:=\left\{T:H^{r}(\partial\Omega)\to L^{2}(\partial\Omega),\quad T=\sum_{j,k=1}^{\infty}t_{jk}b_{jk}^{(r)}:t_{jk}\in\mathbb{R},\sum_{j,k=1}^{\infty}t_{jk}^{2}<\infty\right\}.$
The elements of $\mathbb{H}_{r}$ are the ‘Hilbert–Schmidt’ operators between $H^{r}(\partial\Omega)$ and $L^{2}(\partial\Omega)$. Furthermore, $\mathbb{H}_{r}$ is itself a Hilbert space with inner product
$\langle S,T\rangle_{\mathbb{H}_{r}}:=\sum_{j,k=1}^{\infty}s_{jk}t_{jk}\equiv\sum_{j,k=1}^{\infty}\langle S\phi_{j}^{(r)},\phi_{k}^{(0)}\rangle_{L^{2}(\partial D)}\langle T\phi_{j}^{(r)},\phi_{k}^{(0)}\rangle_{L^{2}(\partial D)}.$
From [1, Lemma 18], we know that $\tilde{\Lambda}_{\gamma}\in\mathbb{H}_{r}$ for any $r$. Consider the set $D_{\Lambda}=\{\tilde{\Lambda}_{\gamma}:\gamma\in\Gamma\}$ of N-t-D operators corresponding to admissible conductivities. We show in Lemma 7.1 that $D_{\Lambda}$ is a compact set of $\mathbb{H}_{r}$.
@@
Formally, the problem is to ‘learn’ the map $G^{\text{true}}:D_{\Lambda}\to\Gamma$. More particularly, we want to design a deep neural network (DeepONet) to approximate the map, $G^{\text{true}}$. For learning operators between infinite dimensional spaces, DeepONets were introduced in [44, 51]. These generalized the (shallow-net) architecture proposed in [11, 12]. Theoretical guarantees for DeepONets were given in [44] where the neural network tries to learn operators between two Hilbert spaces. This theory was further generalized in [40, 42] and an architecture called Fourier Neural Operators was introduced for learning operators between Banach spaces having a certain approximation property. Fairly recently, DeepONets and FNOs have been synthesized as Neural Inverse Operators and they have been used effectively in solving several inverse problems, [55]. With this background in mind, and to set the stage for formulating our learning problem as learning an operator between two Hilbert Spaces, let us introduce a (regular)‘link function’ $\Phi:\mathbb{R}\to[m,\infty)$ such that any conductivity $\gamma\in\Gamma$ is given by $\gamma=\Phi\circ f$ for some $f\in L^{\infty}(\Omega)\cap L^{2}{(\Omega)}:=\tilde{\Gamma}$. An example of such a link function is given in [57, Example 8]. Now we will modify our operator learning problem. For this we first define a map, $\mathcal{G}^{\dagger}:D_{\Lambda}\to\tilde{\Gamma}$ where $\mathcal{G}^{\dagger}=\Phi^{-1}\circ G^{\text{true}}$. Thus knowing $\mathcal{G}^{\dagger}$ completely determines $G^{\text{true}}$. Next we extend this map: Let $\mathcal{G}^{\dagger}:D_{\Lambda}\to L^{\infty}(\Omega)\cap L^{2}(\Omega)$ be the map as above, then there exists an extension $\mathcal{G}^{\dagger}_{\text{ext}}:\mathbb{H}_{r}\to L^{\infty}(\Omega)\cap L^{2}(\Omega)$ such that $\mathcal{G}^{\dagger}_{\text{ext}}|_{D_{\Lambda}}=\mathcal{G}^{\dagger}$. The extension $\mathcal{G}^{\dagger}_{\text{ext}}$ exists by Lemma 7.4. The map $\mathcal{G}^{\dagger}_{\text{ext}}$ will be approximated by a deep neural operator called DeepONet. Essentially, a neural operator is a parametric map between the input space (e.g. space of N-t-D operators) and the output space (e.g. space of admissible conductivities) that can be interpreted as a composition of three maps, an encoder, an approximator and a reconstructor, see [44] and Fig. 1. As such, an upper bound on the error in approximating the true operator by a neural operator can be split into upper bounds on the encoder error, the approximator error and the reconstructor error correspondingly. Now we recall the components of a DeepONet as proposed in [44, 51] and adapt the definitions as applicable to the case of EIT. To that end, we define the following operators while keeping the terminology the same as in [44]:
@@
    (8) $\mathcal{E}_{M}:=\mathcal{E}:D_{\Lambda}\subset\mathbb{H}_{r}\to\mathbb{R}^{M}\times\mathbb{R}^{M};\quad\tilde{\Lambda}_{\gamma}\mapsto\tilde{R}_{M}(\gamma)$
    The encoder can be thought of as ‘sensor evaluations’ in the spirit of [44]. Whenever it is clear from context, we will denote $\mathcal{E}_{M}$ by $\mathcal{E}$. It follows from Lemmas 7.5 and 7.6 that the map $\mathcal{E}$ is continuous and from Lemma 7.4 can be extended continuously on $\mathbb{H}_{r}$. With a slight abuse of notation, we call this extended map $\mathcal{E}$ too.
  • Decoder: While the encoder is a projection of an infinite-dimensional object (i.e. $\tilde{\Lambda}_{\gamma}\in\mathbb{H}_{r}$) into a finite-dimensional space, a decoder lifts back a finite dimensional representation, $R_{M}(\gamma)$ into its corresponding map in the space of Hilbert-Schmidt operators. This lifting is done using the construction given in [21, section 6] and results in an operator ${\Lambda}^{M}_{\gamma}$, see [21, eqn. 6.8]. Thus, we have:
    (9) $\mathcal{D}_{M}:=\mathcal{D}:\mathbb{R}^{M}\times\mathbb{R}^{M}\to\mathbb{H}_{r};\quad\tilde{R}_{M}(\gamma)\mapsto\tilde{\Lambda}^{M}_{\gamma}.$
@@
    (10) $\mathcal{A}:\mathbb{R}^{M}\times\mathbb{R}^{M}\to\mathbb{R}^{P};\quad\tilde{R}_{M}(\gamma)\mapsto\{\mathcal{A}_{k}\}_{k=1}^{P}$
@@
    (11) $\mathbf{\tau}:\mathbb{R}^{d}\to\mathbb{R}^{P+1},\quad y=(y_{1},\cdots,y_{d})\mapsto\{\tau_{k}\}_{k=0}^{P}$
@@
    (12) $\mathcal{R}_{P}:=\mathcal{R}:\mathbb{R}^{P}\to L^{2}(\Omega),\quad\mathcal{R}(\alpha_{k}):=\tau_{0}(y)+\Sigma_{k=1}^{P}\alpha_{k}\tau_{k}(y).$
@@
    (13) $\mathcal{P}_{P}:=\mathcal{P}:L^{2}(\Omega)\to\mathbb{R}^{p},\quad\mathcal{P}(f)=\left(\langle f-\tau_{0},\tau_{1}^{*}\rangle,\cdots,\langle f-\tau_{0},\tau_{P}^{*}\rangle\right)$